OpenStudy (anonymous):

Find the slope and y-intercept for the line x + y = 2 and draw its graph

OpenStudy (fifciol):

determine y so that you'll have form y= mx+b

OpenStudy (anonymous):

y = (-1)x +2 yint (0,2) m = -1

OpenStudy (fifciol):

that's correct, can you plot that function?

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (fifciol):

yes I can

OpenStudy (anonymous):

Determine the correct equation for the line passing through the point (12, –4) and parallel to the line x + 4y = 16.

OpenStudy (anonymous):

my answer is y = -1/4x - 1?

OpenStudy (fifciol):

correct